-2X-5>=-15
-8<=-5+1/3X
5-2X>2X+21
WHAT WOULD THE 7TH NUMBER NEED TO BE IN ORDER TO HAVE A MEAN OF EXACTLY 250 PEOPLE IN ATTENDANCE? 203,256,211,225,276,200

Answer by venugopalramana(3286)

WHAT WOULD THE 7TH NUMBER NEED TO BE IN ORDER TO HAVE A MEAN OF EXACTLY 250 PEOPLE IN ATTENDANCE? 203,256,211,225,276,200
let the 7th.number =x
so mean =sum of all 7 numbers /7=(203+256+211+225+276+200+x)/7=250
(203+256+211+225+276+200+x)=7*250=1750
x=1750-(203+256+211+225+276+200)=1750-1371=379
so the 7th.number is 379
REGARDING INEQUALITIES SEE FOLLOWING EXAMPLES SOLVED BY ME EARLIER AND DO YOUR SELF.IF STILL THERE IS DIFFICULTY COME BACK
Graphs/17660: 2<5-2x/3<7
1 solutions
--------------------------------------------------------------------------------
Answer 8553 by venugopalramana(298) on 2005-10-30 11:25:19 
2<5-2x/3<7..take one by one to solve
2<5-2x/3
2x/3<5-2=3
x<3*3/2
x<9/2
5-2x/3<7
5-7<2x/3
-2<2x/3...or..-2*3/2<1*x...or.....-3<1*x
....so the answer is -3<1*x<9/2
********************************
Graphs/17965: how do you solve eniqualities
1 solutions
--------------------------------------------------------------------------------
Answer 8675 by venugopalramana(298) on 2005-11-01 11:13:46 
the rules of transformations between equalities and inequalities are same but for few exceptions..they are given below..use them to solve the inequalities
...........property..............equality....inequality.......conclusion
...example........................3+2=5.......5<8
1.reverse sides.left to right ...5=3+2......8>5..inequality changes from < to >
2.multiply/divide with -ve number.-5=-(3+2)..-5>-8 same as above.........
3.take reciprocal.................1/(3+2)=1/5..1/5>1/8..same as above..
so take care of these and solve
1example...x+5>8..or x>8-5=3..or x>3
2.3>5+x..or x+5<3..or x<3-5=-2..or x<-2..(exception rule no.1 above)
2.example...-x+5>8..or-x>8-5=3..or -x>3..or x<3(exception rule no2. above)
3.example...1/x>1/8...or x<8....(exception rule no.3 above)
SO BE CAREFULL IN THESE OPERATIONS..PARTICULARLY MULTIPLICATION/DIVISION WITH UNKNOWNS..YOU HAVE TO BE SURE WHETHER IT IS A +VE NUMBER OR _VE NUMBER